OpenStudy (anonymous):

can you simplify 3 radical 128 step by step

OpenStudy (aum):

\[128 = 2 \times 2 \times 2 \times 2 \times 2 \times 2 \times 2 = \large 2^7 = 2^3 * 2^3 * 2 \\ \large \sqrt[3]{128} = (128)^{1/3} = (2^3 * 2^3 * 2)^{1/3} = (2^3)^{1/3} * (2^3)^{1/3} * (2)^{1/3} = \\ \large 2^1 * 2^1 * (2)^{1/3} = 4\sqrt[3]{2} \]